OpenStudy (anonymous):

Graph the circle with center at (3, -2), which also passes through the point (0,2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (gamer56):

Hi,id love to help but I am not doing it for you. What are you confused about? What are you stuck on?

OpenStudy (anonymous):

Try thinking about what you have here. The centre, a point on the circle. What can you find using that information?

OpenStudy (anonymous):

u can find distance right

OpenStudy (anonymous):

@JFlaus

OpenStudy (anonymous):

Exactly, and what's another way of saying the distance between the centre and a point on the circle?

OpenStudy (anonymous):

center?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

the center of the circle

OpenStudy (anonymous):

Yes, we have the centre of the circle and a point on the circle. So from that we can get the radius right? :)

OpenStudy (anonymous):

mmhm so i need to do the distance formula first

OpenStudy (anonymous):

@JFlaus

OpenStudy (anonymous):

Yes, ultimately our goal is to get the equation of the circle right? And for that we can use the equation formula which requires us to have things like radius :) So were workig them out from what we've been given :)

OpenStudy (anonymous):

ok ill do the distance formula and find that out ill tag u when i get it

OpenStudy (anonymous):

@JFlaus it says 5 is the distance

OpenStudy (anonymous):

Sounds good, so tell me all the stuff we know about this circle.

OpenStudy (anonymous):

we know both points and the center

OpenStudy (anonymous):

Good, so tell me what the formula for equation of a circle is.

OpenStudy (anonymous):

(x−h)2+(y−k)2=

OpenStudy (anonymous):

@JFlaus

OpenStudy (anonymous):

= r^2

OpenStudy (anonymous):

So we take the formula (x-h)^2+(y-k)^2=r^2 and we sub in what we have. So what do h and k refer to? :)

OpenStudy (anonymous):

ik h is height i forgot what k is

OpenStudy (anonymous):

Not quite. H would be height in another question, but in this formula you take the centre point as (h,k) :)

OpenStudy (anonymous):

So if the centre is (3, -2) then h = 3 and k = -2 :)

OpenStudy (anonymous):

ok so now i just plug it in? @JFlaus

OpenStudy (anonymous):

Yup :) And thatll give you the equation of the circle :)

OpenStudy (anonymous):

(x−3)2+(y−-2)2= R^2 ?

OpenStudy (anonymous):

Now solve it out :)

OpenStudy (anonymous):

wait isnt R=5

OpenStudy (anonymous):

Yeah r was 5 so we sub it in and square it :)

OpenStudy (anonymous):

ok so whats the formula completly

OpenStudy (anonymous):

@JFlaus

OpenStudy (anonymous):

The formula is (x-h)^2+(y-k)^2=r^2 :)

OpenStudy (anonymous):

ok and do i just graph that

OpenStudy (anonymous):

@JFlaus sorry last question ^

OpenStudy (anonymous):

( X-3)^2 + (Y +2 )^2= 5^2

OpenStudy (anonymous):

No, when you simplify that out you get the equation of the circle and thats your answer isnt it? :P

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

@JFlaus

OpenStudy (anonymous):

The question asks you to find the equation of the circle right? :)

OpenStudy (anonymous):

yes how would u simplfy this

OpenStudy (anonymous):

You take the formula and you sub in for h and k and r and then you simplify it :)